OpenStudy (anonymous):

solving an equation graphically 2x+y=-8,x=5y-5

OpenStudy (klimenkov):



OpenStudy (anonymous):

but on the math i got stuck and can't figure out how it came to the conculsion

OpenStudy (klimenkov):

Plot it very accurate and find the coordinates of the point of intersection.

OpenStudy (anonymous):

i mean on the math i got it to this point ( one sec) 2x+y=-8 y=2x-8 then i get stuck

OpenStudy (klimenkov):

Plot 2 lines: \(y=-2x-8, \; y=\frac x5-1\) Can you do this?

OpenStudy (anonymous):

no this is my first day of doing these problems trying to perfect the math

OpenStudy (klimenkov):

http://www.wolframalpha.com/input/?i=Plot [{-2x-8%2Cx%2F5%2B1}%2C{x%2C-7%2C7}]

OpenStudy (klimenkov):

Copy all the link and paste into the browser.

OpenStudy (anonymous):

its taking me to a deadlink

OpenStudy (klimenkov):

http://www.wolframalpha.com/ Is this a deadlink too?

OpenStudy (anonymous):

no

OpenStudy (klimenkov):

type in the box: Plot[{-2x-8,x/5+1},{x,-7,7}]

OpenStudy (anonymous):

are you doing this on a calculator?

OpenStudy (klimenkov):

Actually not. But it is easy to plot lines in this way. My picture that you saw higher has lower quality that this one.

OpenStudy (anonymous):

the first one i got (-4,8) and the second one i can't get past y=5-x/5

OpenStudy (klimenkov):

Solution is \((-\frac{45}{11},\frac2{11})\).

OpenStudy (anonymous):

k but how did you get there cause i don't even know plz and thanxs

OpenStudy (klimenkov):

1. If you want to solve it graphically you have to plot both of lines at the same time. And the point, where this lines intersect will be a solution. This is not an accurate method, because you can draw lines perfectly well. 2. If you want an accurate solution, just put an expression for x from the second equation to the first one. you will get: \(2(5y-5)+y=-8\) \(10y-10+y=-8\) \(11y=2\;\Rightarrow\; y=\frac2{11}\) To find an x put the value of y into one of the equations: \(2x+\frac2{11}=-8 \; \Rightarrow \; x=-\frac{45}{11}\)

OpenStudy (anonymous):

thanxs